OpenStudy (anonymous):

How to find the area of a circle just given the circumference? The circumference is 64pi

jimthompson5910 (jim_thompson5910):

Circumference = 2*pi*radius C = 2*pi*r 64pi = 2*pi*r solve for r to get the radius

jimthompson5910 (jim_thompson5910):

once you have the radius, you can use it to find the area using the formula A = pi*r^2

OpenStudy (anonymous):

How do I find the radius?

jimthompson5910 (jim_thompson5910):

what do you get when you solve 64pi = 2*pi*r for r?

OpenStudy (anonymous):

64/3.14=20.382/2=10.1910

jimthompson5910 (jim_thompson5910):

64pi = 2*pi*r 2*pi*r = 64pi r = 64pi/(2*pi) r = 64/2 r = 32

OpenStudy (anonymous):

I thought it was divided by pi? Ok let me plug that into the other equatin you gave me

jimthompson5910 (jim_thompson5910):

64pi divided by 2pi = 32 because the two "pi" terms cancel and you're left with 64/2 = 32

OpenStudy (anonymous):

I keep getting 200.96 but thats not right

jimthompson5910 (jim_thompson5910):

A = pi*r^2 A = 3.14*32^2 A = 3215.36

OpenStudy (anonymous):

Thats not one of the answers given.

jimthompson5910 (jim_thompson5910):

what are your answer choices?

OpenStudy (anonymous):

I got 1023.

OpenStudy (anonymous):

I did A = 3.14 x (r x r) A = 3.14 x 326.045 A = 1023

jimthompson5910 (jim_thompson5910):

how did you get 326.045 ?

jimthompson5910 (jim_thompson5910):

btw, what are your answer choices?

OpenStudy (anonymous):

1. 128pi in^2 2. 256 pi in^2 3. 1024 pi in^2 4. 16 pi in^2

jimthompson5910 (jim_thompson5910):

oh you just leave the pi alone

jimthompson5910 (jim_thompson5910):

A = pi*r^2 A = pi*(32)^2 A = pi*(1024) A = 1024pi

OpenStudy (anonymous):

Thanks

jimthompson5910 (jim_thompson5910):

np